# mCLM: A Function-Infused and Synthesis-Friendly Modular Chemical Language Model

### v1

@@ Abstract
Despite their ability to understand chemical knowledge and accurately generate sequential representations, large language models (LLMs) remain limited in their capacity to propose novel molecules with drug-like properties. In addition, the molecules that LLMs propose can often be challenging to make in the lab. To more effectively enable the discovery of functional small molecules, LLMs need to learn a molecular language. However, LLMs are currently limited by encoding molecules from atoms. In this paper, we argue that just like tokenizing texts into (sub-)word tokens instead of characters, molecules should be decomposed and reassembled at the level of functional building blocks, i.e., parts of molecules that bring unique functions and serve as effective building blocks for real-world automated laboratory synthesis. This motivates us to propose mCLM, a modular Chemical-Language Model tokenizing molecules into building blocks and learning a bilingual language model of both natural language descriptions of functions and molecule building blocks. By reasoning on such functional building blocks, mCLM guarantees to generate efficiently synthesizable molecules thanks to recent progress in block-based chemistry, while also improving the functions of molecules in a principled manner. In experiments on 430 FDA-approved drugs, we find mCLM capable of significantly improving 5 out of 6 chemical functions critical to determining drug potentials. More importantly, mCLM can reason on multiple functions and improve the FDA-rejected drugs (“fallen angels”) over multiple iterations to greatly improve their shortcomings.  [^1] All codes, data, and models will be released publicly upon publication.
@@ 1 Introduction
Figure 1: mCLM adopts a modular chemical vocabulary, which uses molecular modules as tokens together with natural language tokens. Note that we also refer to these modules as “building blocks”. Compared with using natural language names, SMILES strings, or holistic embeddings for whole molecules, this level of tokenization aligns better with molecular function groups. It allows more efficient molecular structure modification and synthesis.
@@
Small molecules—the class of chemical matter primarily built from carbon atoms bonded together—can perform a wide range of important functions in human society. These include essentials like promoting health by acting as medicines [125, 20, 81, 90, 110], converting energy by functioning as key components in solar cells [69, 65, 48, 49, 79], and achieving sustainability by serving as inherently recyclable products. These functions also include many nice-to-haves that drive substantial economic growth, including acting as colorants, flavorings, perfumes, cosmetics, lotions, sunscreens, coatings, quantum dots, molecular computers, chemical biology probes, insect repellants, etc.
@@
A major challenge with small molecules, however, is that the traditional approach for small molecule synthesis is highly artisanal, slow, and expensive. It considers all possible arrangements of atoms and centers on creating for each targeted molecule a unique process for its assembly that leverages a menu of thousands of different reactions each run under thousands of possible conditions and using millions of possible starting materials. This approach imposes several limitations: (1) It is unfriendly to automation – machines are not good at doing thousands of different things thousands of different ways from a million different starting points. (2) It leads to an undemocratized landscape in drug discovery. With development costs averaging around \$1.3 billion per drug, only economically advanced countries can afford to invest in such high-risk research  [41]. Moreover, participation in the process of molecular innovation currently requires access to highly trained experts in chemical synthesis. (3) Most importantly, there are many instances of known commercial drugs or materials that have well-documented limitations that have remained unaddressed. For example, Imatinib is an important anticancer drug that works well in most parts of the body, but it only poorly penetrates the blood-brain barrier [87, 77, 14, 36]. This means it may effectively treat cancer at its primary site, yet fail to prevent fatal brain metastasis. In such cases, more blood-brain-barrier penetrant molecules that retain all of the other favorable properties of the current drug are desired, but making such selective modifications in functional properties can be very challenging. As another example in the materials domain, organic photovoltaic (OPV) molecules are expected to be more economical and environment-friendly alternatives to current solar cells. However, current commercialized OPV devices either achieve less than 10% energy conversion efficiency—significantly lower than traditional silicon solar cells—or have stability far short of 10 years  [82]. Overcoming these types of drawbacks of current molecules remains a challenging problem in the domain.
An alternative block-based chemistry approach for small molecule synthesis has recently emerged [28, 107, 50, 45, 92, 8, 4, 100, 84], and we recognized that this approach provides an opportunity to reimagine the intersection of AI and small molecule science. Block-based chemistry is chemistry that machines can do. It iteratively assembles small molecules from prefabricated building blocks using chemistry that is simple and general and thus readily automated. Akin to automated DNA, RNA, and peptide synthesis platforms, a major strength of this block-based approach is that it can access billions of novel small molecules with high degrees of functional potential using only a few automation-friendly reactions and a bounded set of pre-fabricated function-infused building blocks.
@@
We specifically recognized that this block-based approach provided an opportunity to create a new modular language for chemistry. The idea that molecules and their synthesis may be best understood from the perspective of a “chemical language” dates back to 2014 [11], where structural fragments and functional groups play the roles of “chemical words”. This view is supported by multiple observations aligned with natural language: (1) they can be decomposed and reassembled, (2) they exhibit ambiguity—the same building block can perform vastly different functions depending on the chemical context, and (3) they possess significant diversity, as many different structures can lead to the same function. This linguistic parallel suggests the potential to train a large language model specifically for molecules by linearizing their structures into modular sequences.
@@
A common representation for such an approach is the Simplified Molecular Input Line Entry System (SMILES) [104], where atoms are denoted by one- or two-character symbols (e.g., C for carbon, Br for bromine, and F for fluorine), rings are represented with numbers, and branches are indicated using parentheses. However, such atom-level tokenization strategies (e.g., SMILES [104] or SELFIES [43]) resemble character-level natural language models, which struggle to generalize effectively. Unlike proteins, which have a fixed vocabulary of 20 amino acids, small molecules exhibit an open vocabulary when each atom is treated as a token, as illustrated in Figure 1. It also causes severe restrictions in practicality because many of the new structures proposed by LLMs based on atom-level tokenization are not practically synthesizable in the laboratory. This approach has created a major gap between what is now possible in silico and what is possible in the physical world.
@@
Furthermore, the SMILES representation can obscure critical structure information: two atoms that are direct neighbors in the molecular graph may be distantly separated in the SMILES string. Even with recent efforts to align SMILEs and natural language description [17, 73, 2], integrate chemical properties and functional groups [71] into SMILES, incorporate 3D geometric information [25, 54], and employ graph neural networks to capture molecular graphs and chemical reaction contexts [97], these representations still fail to encapsulate functional knowledge that is often described only in natural language literature because the inherent properties and functions of molecules are hidden in their structure, composition, and interaction.
@@
Against this backdrop, we argue that there is a correctable fundamental mismatch between the way LLMs work and the way chemists traditionally synthesize and study small molecules. Increasing the “tokenization granularity” down to the corresponding letters is not helpful, and in fact makes the extraction of meaning much more complex and causes generative AI to hallucinate words that don’t exist. To bridge this gap, in this paper, we aim to teach computers to speak two complementary languages: one that represents molecular building blocks (i.e., subgraph structures) indicative of specific functions and compatible with automated modular assembly, and another that describes these functions in natural language (Figure 2). Unlike existing approaches that add such knowledge as a post hoc step, we developed a function- and synthesis-aware modular chemical language model (mCLM). Inspired by bilingual speakers who frequently “code-switch” (naturally and often switch between their two languages within the same message), we propose a novel neural encoder that integrates molecular structure and natural language. mCLM incorporates both function- and synthesis-related knowledge into the small molecule tokenization process a priori.
First, we tokenize small molecules at the level of building blocks (graph substructures) that are better able to predict function and are, by design, fully compatible with automated modular small molecule synthesis. We use graph neural networks to encode each building block. We then extract natural language sentences from the literature that describe the molecule’s functions and chemical reactions and synthesis constraints of various molecules, and seamlessly insert the encoded building blocks alongside the corresponding entity names, as illustrated in Figure 2. During the inference stage, mCLM enables these functional modules to be predictably and automatically assembled into new molecules with desired functions.
@@
mCLM has multiple potential advantages of encoding and generating molecules at a modular level:
@@
    Synthesis efficiency: synthesis can be faster and more broadly accessible because the process is, by design, simple, iterative, general, and machine-friendly. This can enable rapid iterative drug and material development and even automated lab experiments.
@@
    Alignment with language models: resembling the mechanism of natural language tokenization, molecular modularization provides a more natural interface to align with word representations.
@@
    Reasoning: leveraging LLMs’ instruction following capability and wide-scale pretraining, mCLM is able to iteratively refine molecules based on knowledge of molecular functions.
@@ 2 The Modular Chemical Language Model
2.1 Overview
Figure 2: An overview of the mCLM. It is a multimodal chemical-language model that tokenizes molecules into building blocks. After being trained on datasets consisting of properties, functions and synthesis data, the mCLM can conduct critical chemical reasoning through an iterative process to refine molecular design.
Small molecules pose unique challenges to tokenize and encode due to their perceived open-ended structural diversity. This has caused the field to consider them differently from natural language where well-defined vocabularies exist. In natural language modeling, tokenization identifies common substrings (such as words and subwords), which carry richer semantic information than sequences of characters. We note that most small molecules that occur in nature are similarly composed primarily of connected building blocks  [45, 92]. There is likewise a high degree of inherent modularity in many medicines and materials  [23, 22, 3, 5, 45, 92, 95]. We propose mCLM, a multimodal model that jointly encodes and understands natural language and molecules based on synthesis-friendly building blocks instead of atoms. In Section 2.2, we introduce the concept of molecular building blocks as a chemical “vocabulary” in addition to the natural language vocabulary we use. Section 2.3 then describes the tokenization process to obtain the library of building blocks. Then in Section 2.4 we describe how the mCLM trains and runs on the token sequences. Finally, we introduce the reasoning mechanism of mCLM which refines molecule design over multiple iterations in Section 2.5.
@@ 2.2 A Function-Infused and Synthesis-Friendly Vocabulary
In this work, we propose to leverage a chemical vocabulary of synthesis-friendly building blocks. This approach guarantees capacity for automated iterative assembly [51]. The blocks can be chemically connected using predefined synthesis rules in a short period. In particular, we propose a chemically meaningful vocabulary, $V$. Briefly, akin to language models developed for peptides/proteins, small molecules can be assembled automatically from accessible building blocks. These building blocks are often highly associated with chemical functions, such as binding to protein targets, modulating enzyme activity, or affecting involved metabolic processes. This will enable rapid and iterative proposal of new small molecules, automated synthesis of those small molecules, and generation of the corresponding functional data on demand. Notably, in contrast to SMILES strings which break molecule structure during graph linearization (e.g., separating physically adjacent subgraphs such as the two carbons in molecule C(N)C), our representation is linear in the physical world.
2.3 Tokenization
@@
To generate molecules that are not only valid but also automatically synthesizable, we tokenize molecular inputs into molecular “building blocks” (i.e., subgraphs) that align with known chemical synthesis procedures. During training, we combine two tokenization strategies to balance coverage and synthetic feasibility: a synthesis-guaranteed tokenizer and a rule-based tokenizer (Figure 3). The synthesis-guaranteed tokenizer disconnects the molecule only at bonds that can be formed by a predetermined small set of reactions that can be performed in an automated manner. Specifically, it only permits bond disconnections that correspond to reactions compatible with state-of-the-art automated synthesis platforms. These three bonds are: amide coupling, Suzuki-Miyaura coupling, and Buchwald-Hartwig coupling (see Figure 3 tokenization rules) [93] . Unlike traditional retrosynthesis tools that prioritize maximum coverage or human-designed heuristics, this tokenizer is guided by the operational constraints of block-based chemistry. It also ensures that resulting blocks are free from functional group conflicts that would interfere with downstream synthesis. Please see Appendix D.1 for more information. This conservative approach leads to a set of molecular tokens that are guaranteed to work under defined synthesis protocols, enabling seamless transition from model-generated output to real-world synthesis. When the synthesis-guaranteed tokenizer cannot fully cover a molecule—due to chemical incompatibilities or reaction constraints—we fall back on a rule-based tokenizer to ensure coverage of a larger variety of molecules. The rule-based tokenizer also breaks molecules along the same automated machine-friendly bonds as the above tokenizer. However, no other rules are applied beyond specifying a minimum size for blocks. This rule-based tokenizer is used only during training to support learning on diverse molecular structures while maintaining consistency with synthesis logic.

### v2

# mCLM: A Modular Chemical Language Model that Generates Functional and Makeable Molecules

@@ Abstract
Despite their ability to understand chemical knowledge, large language models (LLMs) remain limited in their capacity to propose novel molecules with desired functions (e.g., drug-like properties). In addition, the molecules that LLMs propose can often be challenging to make, and are almost never compatible with automated synthesis approaches. To better enable the discovery of functional small molecules, LLMs need to learn a new molecular language that is more effective in predicting properties and inherently synced with automated synthesis technology. Current molecule LLMs are limited by representing molecules based on atoms. In this paper, we argue that just like tokenizing texts into meaning-bearing (sub-)word tokens instead of characters, molecules should be tokenized at the level of functional building blocks, i.e., parts of molecules that bring unique functions and serve as effective building blocks for real-world automated laboratory synthesis. This motivates us to propose mCLM, a modular Chemical-Language Model that comprises a bilingual language model that understands both natural language descriptions of functions and molecular blocks. mCLM front-loads synthesizability considerations while improving the predicted functions of molecules in a principled manner. Experiments on 430 FDA-approved drugs showed that mCLM is capable of significantly improving chemical functions critical to determining drug potentials. mCLM, with only 3B parameters, also achieves improvements in synthetic accessibility relative to 7 other leading generative AI methods including GPT-5. When tested on 122 out-of-distribution medicines using only building blocks/tokens that are compatible with automated modular synthesis, mCLM outperforms all baselines in property scores and synthetic accessibility. mCLM can also reason on multiple functions and iteratively self-improve to rescue drug candidates that failed late in clinical trials (“fallen angels”).  [^1] All codes, data, and models will be released publicly upon publication.
@@ 1 Introduction
Figure 1: mCLM adopts a modular chemical vocabulary, which uses synthesis robot-friendly molecular building blocks as tokens together with natural language tokens. Compared with using natural language names, SMILES strings, or holistic embeddings for whole molecules, this building block level of tokenization better enables the prediction of compounds with improved properties, and guarantees automated synthesizability a priori, thus building a direct link between the digital and physical worlds. This approach stands to substantially enable AI-guided discovery of new small molecules with targeted functions.
@@
Small molecules—the class of chemical matter primarily built from carbon atoms bonded together—can perform a wide range of important functions in human society (Zhang et al., 2025). These include essentials like promoting health by acting as medicines (Zheng et al., 2024; Edwards et al., 2024b; Singhal et al., 2023; Thirunavukarasu et al., 2023; Xiao et al., 2024a), converting energy by functioning as key components in solar cells (Nguyen et al., 2024b; Lv et al., 2021; Li et al., 2023b; 2024b; Si et al., 2024), and achieving sustainability by serving as inherently recyclable products. These functions also include many nice-to-haves that drive substantial economic growth, including colorants, flavorings, perfumes, cosmetics, coatings, quantum dots and insect repellants.
@@
The traditional approach for small molecule synthesis is highly artisanal, and thus slow and expensive: (1) It is unfriendly to automation: machines are not good at performing thousands of reaction types, with each run under thousands of possible conditions and using millions of possible starting materials. (2) It leads to an undemocratized landscape in drug discovery. With development costs averaging around \$1.3 billion per drug, only economically advanced countries can afford to invest in such high-risk research  (Kneller, 2010). It also excludes potential breakthroughs from the developing world and inhibits research that could prevent diseases prevalent there. Moreover, participation in the process of molecular innovation currently requires access to highly trained experts in chemical synthesis. As a result, there are no effective drugs for many diseases such as Parkinson and Amyotrophic Lateral Sclerosis. (3) Most importantly, there are many instances of known commercial drugs or materials that have well-documented limitations that have remained unaddressed. For example, Imatinib is an important anticancer drug that works well in most parts of the body, but it only poorly penetrates the blood-brain barrier (Takayama et al., 2002; Senior, 2003; Coutre et al., 2004; Isobe et al., 2009). This means it may effectively treat cancer at its primary site, yet fail to prevent fatal brain metastasis. In such cases, more blood-brain-barrier penetrant molecules that retain all of the other favorable properties of the current drug are desired, but making such selective modifications in functional properties can be very challenging. As another example in the materials domain, 350+ million people in Asia and the Pacific have only limited access to electricity, and 150 million people still have no access at all. Organic photovoltaic (OPV) molecules are expected to be more economical and environment-friendly alternatives to current solar cells. However, current commercialized OPV devices either achieve less than 10% energy conversion efficiency—significantly lower than traditional silicon solar cells—or have stability far short of 10 years  (Solak and Irmak, 2023).
@@
An alternative block chemistry approach for small molecule synthesis has recently emerged (Gillis and Burke, 2007; Woerly et al., 2014; Li et al., 2015a; Lehmann et al., 2018; Trobe and Burke, 2018; Blair et al., 2022a; Angello et al., 2022; Wang et al., 2024; Strieth-Kalthoff et al., 2024a). Block chemistry is iterative carbon-carbon bond-forming chemistry that machines can do. It iteratively assembles small molecules from prefabricated building blocks using chemistry that is simple and general and thus readily automated. Akin to automated DNA, RNA, and peptide synthesis platforms, a major strength of this block-based approach is that it can access billions of novel small molecules with high degrees of functional potential using only a few automation-friendly reactions and a bounded set of pre-fabricated function-infused building blocks. We specifically recognized that this block-based approach provided an opportunity to create a new modular language for chemistry. The idea that molecules and their synthesis may be best understood from the perspective of a “chemical language” dates back to 2014 (Cadeddu et al., 2014), where structural fragments and functional groups play the roles of “chemical words”. This view is supported by multiple observations aligned with natural language: (1) they can be decomposed and reassembled, (2) they exhibit ambiguity—the same building block can perform vastly different functions depending on the chemical context, and (3) they possess significant diversity, as many different structures can lead to the same function. This linguistic parallel suggests the potential to train a large language model specifically for molecules by linearizing their structures into modular sequences.
@@
A common representation for such an approach is the Simplified Molecular Input Line Entry System (SMILES; Weininger, 1988), where atoms are denoted by one- or two-character symbols (e.g., C for carbon, Br for bromine, and F for fluorine), rings are represented with numbers, and branches are indicated using parentheses. However, such atom-level tokenization strategies (e.g., SMILES (Weininger, 1988) or SELFIES (Krenn et al., 2020)) resemble character-level natural language models, which struggle to generalize effectively. Unlike proteins, which have a fixed vocabulary of 20 amino acids, small molecules exhibit an open vocabulary when each atom is treated as a token, as illustrated in Figure 1. It also causes severe restrictions in practicality because many of the new structures proposed by LLMs based on atom-level tokenization are not practically synthesizable in the laboratory. This approach has created a major gap between what is now possible in silico and what is possible in the physical world.
@@
Furthermore, the SMILES representation can obscure critical structure information: two atoms that are direct neighbors in the molecular graph may be distantly separated in the SMILES string. Even with recent efforts to align SMILES and natural language description (Edwards et al., 2022a; Pei et al., 2024; Ahmad et al., 2022), integrate chemical properties and functional groups (Nguyen et al., 2025) into SMILES, incorporate 3D geometric information (Fu et al., 2025; Li et al., 2025a), and employ graph neural networks to capture molecular graphs and chemical reaction contexts (Wang et al., 2022a), these representations still fail to encapsulate functional knowledge that is often described only in natural language literature because the inherent properties and functions of molecules are hidden in their structure, composition, and interaction.
Therefore, our goal is not to make all drug-like molecules, rather, our goal is to figure out how to make the right ones, better, faster, stronger. Unlike machines, human scientists are inherently “multilingual,” seamlessly navigating diverse modalities—from natural language and scientific figures in literature to complex scientific data such as molecular structures in knowledge bases. In contrast, most prior work on molecule discovery trains large language models on a single modality. Moreover, Human scientists “think before they talk,” grounding their reasoning in deeply reflective and deliberate reflection and critical evaluation to generate new hypotheses. Current models lack this critical thinking capacity, limiting their ability to contribute meaningfully to discovery. In particular, the human body is a highly complex, interconnected system, and drug discovery is essentially a multi-objective optimization problem. It involves balancing factors such as drug absorption, first-pass metabolism, bioavailability, distribution, protein binding, and clearance. However, improving one property often comes at the expense of another. Many promising drug candidates have failed in the final stages of FDA approval due to an unacceptably high risk of drug-induced liver injury.
@@
Against this backdrop, we argue that there is a correctable fundamental mismatch between the way LLMs work and the way chemists traditionally synthesize and study small molecules. Reducing tokenization granularity to the level of individual letters is counterproductive, as it complicates meaning extraction and increases the likelihood of generative AI hallucinate words that don’t exist. Analogous limitations are inherently linked to atom-based tokenization. And as a result, unfortunately, much of the generative AI research for scientific discovery within the computer science community does not extend to hypothesis verification in the physical world. To bridge this gap, in this paper, we propose a novel LLM by drawing inspiration from the scientific discovery process itself. We aim to develop a science-inspired large language model that follow three principles: (1) “Observe” - acquire, represent and integrate knowledge from multiple data modalities; (2) “Think” – think critically to generate hypotheses; and (3) “Propose” – verify hypotheses through the Physical World. We aim to teach computers to speak two complementary languages: one that represents molecular building blocks (i.e., subgraph structures) indicative of specific functions and compatible with automated modular assembly, and another that describes these functions in natural language (Figure 2). Unlike existing approaches that add such knowledge as a post hoc step, we develop a function- and synthesis-aware modular chemical language model (mCLM). Inspired by bilingual speakers who frequently “code-switch” (naturally and often switch between their two languages within the same message; Poplack, 2013), we propose a novel neural encoder that integrates molecular structure and natural language. mCLM incorporates both function- and synthesis-related knowledge into the small molecule tokenization process a priori. First, we tokenize small molecules at the level of building blocks (graph substructures) that are able to predict function and are, by design, flexible, multi-scale, fully compatible with automated modular small molecule synthesis. We use graph neural networks to encode each building block. We then extract natural language sentences from the literature that describe the molecule’s functions and chemical reactions and synthesis constraints of various molecules, and seamlessly insert the encoded building blocks alongside the corresponding entity names to form the training data, as illustrated in Figure 2. By reasoning on such functional building blocks, mCLM guarantees to generate efficiently synthesizable molecules thanks to recent progress in block-based chemistry, while also improving the functions of molecules in a principled manner. During model inference, mCLM enables these functional modules to be predictably and automatically assembled into molecules with desired functions.
@@
Compared to previous work, mCLM has multiple potential advantages of encoding and generating molecules at a modular level:
@@
    Synthesis efficiency: proposed molecules can be faster and more broadly makeable because the process is, by design, simple, iterative, general, and machine-friendly. This can enable rapid iterative drug and material development via seamless integration with automated lab experiments.
@@
    Multimodal understanding: resembling the mechanism of natural language tokenization, molecular modularization provides a more natural interface to align with word representations to form a more powerful multimodal representation.
@@
    Deliberate reasoning: leveraging LLMs’ instruction following capability and wide-scale pretraining, mCLM is able to iteratively refine molecules based on knowledge of molecular functions.
@@ 2 The Modular Chemical Language Model
Figure 2: An overview of the mCLM. It is a multimodal chemical-language model that tokenizes molecules into synthesis robot-friendly building blocks, thus creating a direct link between the digital and physical worlds. After being trained on datasets consisting of properties, functions and synthesis data, the mCLM can conduct critical chemical reasoning through an iterative refinement process.
In natural language modeling, tokenization identifies common substrings (such as words and subwords), which carry richer semantic information than sequences of characters. Similarly in nature, most small molecules are composed primarily of connected building blocks  (Lehmann et al., 2018; Trobe and Burke, 2018). There is likewise a high degree of inherent modularity in many medicines and materials  (Ertl and Schuhmann, 2019; Arkan et al., 2020; Andrews et al., 1984; Vitaku et al., 2014).To leverage the phenomena in chemistry and borrow the spirit from natural language modeling, we propose mCLM, a multimodal model that jointly encodes and understands natural language and molecules based on synthesis-friendly building blocks instead of atoms. In Section 2.1, we introduce the concept of molecular building blocks as a chemical “vocabulary” and describe the tokenization process to obtain the library of building blocks. Then in Section 2.2 we describe the mCLM architecture and its training. Finally, we introduce the reasoning mechanism of mCLM which refines molecule design over multiple iterations in Section 2.3.
@@ 2.1 A Function-Infused and Synthesis-Friendly Vocabulary
In this work, we propose to leverage a chemical vocabulary $V$ of synthesis-friendly building blocks. This approach guarantees capacity for automated iterative assembly first reported in 2015 (Li et al., 2015b) and, since then, demonstrated in multiple experimental campaigns including Strieth-Kalthoff et al. (2024b) and Angello et al. (2024). The blocks can be chemically connected using predefined synthesis rules in a short period. Briefly, akin to language models developed for peptides/proteins, small molecules can be assembled automatically from makeable building blocks. These building blocks are often highly associated with chemical functions, such as binding to protein targets, modulating enzyme activity, or affecting involved metabolic processes. This will enable rapid and iterative proposal of new small molecules, automated synthesis of those small molecules, and generation of the corresponding functional data on demand. In contrast to SMILES strings which break molecule structure during graph linearization (e.g., separating physically adjacent subgraphs such as the two carbons in molecule C(N)C), our representation is linear in the physical world.
@@
During dataset construction, we combine two tokenization strategies to balance coverage and synthetic feasibility: a synthesis-guaranteed tokenizer and a rule-based tokenizer (Figure 3). The synthesis-guaranteed tokenizer disconnects the molecule only at bonds that can be formed by a predetermined small set of reactions that can be performed in an automated manner. Specifically, it only permits bond disconnections that correspond to reactions compatible with state-of-the-art automated synthesis platforms. These three bonds are: amide coupling, Suzuki-Miyaura coupling, and Buchwald-Hartwig coupling (see Figure 3 tokenization rules; Tyrikos-Ergas et al., 2025). The disconnection rules are adapted from well-established principles in block-based and automated chemistry (Blair et al., 2022b; Li et al., 2015b). Qualitative examples of tokenization are included in the appendix F, and the full rules and tokenizer code will be released with documentation. Importantly, the tokenizer is modular: as new automated reactions are validated in practice, their corresponding disconnection rules can be readily incorporated, expanding the makeable chemical space without requiring fundamental architectural changes. Unlike traditional retrosynthesis tools that prioritize maximum coverage or human-designed heuristics, this tokenizer is guided by the operational constraints of block chemistry. It also ensures that resulting blocks are free from functional group conflicts that would interfere with downstream synthesis. This conservative approach leads to a set of molecular tokens that are guaranteed to work under defined synthesis protocols, enabling seamless transition from model-generated output to real-world synthesis. When the synthesis-guaranteed tokenizer cannot fully cover a molecule—due to chemical incompatibilities or reaction constraints—we fall back on a rule-based tokenizer to ensure coverage of a larger variety of molecules. The rule-based tokenizer also breaks molecules along the same automated machine-friendly bonds as the above tokenizer. However, no other rules are applied beyond specifying a minimum size for blocks. This rule-based tokenizer is used only during training to support learning on diverse molecular structures while maintaining consistency with synthesis logic. We note that the vocabulary growth is analogous to the shift from character- level to subword-level tokenization in natural language, where larger but semantically meaningful tokens significantly improved performance.